OpenStudy (thecalchater):

6 calc questions from homework problems? Area and Definite Integrals?

OpenStudy (thecalchater):

here are the first 3

OpenStudy (thecalchater):

for the first one i got d

OpenStudy (freckles):

\[\sum_{k=3}^5 f(k)=f(3)+f(4)+f(5)\]

OpenStudy (freckles):

the sum of three fractions in the first one should give you d

OpenStudy (thecalchater):

yep ok number 2?

OpenStudy (freckles):

in the next one find the area of each rectangle...then add like look at the first graph, find the area of each rectangle in the first graph... then just add those rectangles then do the same for the second graph

OpenStudy (thecalchater):

ok give me a second

OpenStudy (thecalchater):

I got C. What did you get?

OpenStudy (freckles):

how did you get C

OpenStudy (thecalchater):

here are the other 3. For 1. D 2. 3. 4. 5. B 6. A

OpenStudy (freckles):

did you find the base length of each rectangle?

OpenStudy (thecalchater):

I tried to but I'm kind of confused.

OpenStudy (dan815):

f(ci) is the height

OpenStudy (freckles):

the area lies between x=1 and x=2 and the length from x=1 to x=2 is 1 and you want to divide 1 into 5 parts which means the base length is 1/5 for each rectangle

OpenStudy (thecalchater):

ok so the second one is 1/5 also but the heights are different.

OpenStudy (freckles):

right you will have to use the function given to find the the heights

OpenStudy (thecalchater):

so plug in 1/5 into the equation?

OpenStudy (freckles):

you will need to evaluate the function at each left endpoint of the sub-intervals and you will need to evaluate the function at each right endpoint of the sub-intervals

OpenStudy (thecalchater):

ok so how do i do that?

OpenStudy (freckles):

the function given is f(x)=1/x

OpenStudy (thecalchater):

right

OpenStudy (freckles):

use the length of each sub-interval to determine the left and right endpoints

OpenStudy (thecalchater):

so about (1,1) and (2,0.6) ?

OpenStudy (freckles):

for example the first sub-interval is [1,(1+1/5)] or [1,6/5] the left endpoint here is x=1 the right endpoint here is x=6/5 you have 4 more sub-intervals

OpenStudy (thecalchater):

how do i find them.... sorry i am new to this.

OpenStudy (freckles):

the length of each sub-interval is 1/5....we found this number earlier

OpenStudy (thecalchater):

ok

OpenStudy (freckles):

do you know how to compute 6/5+1/5 this will give you the next right endpoint

OpenStudy (thecalchater):

7/5 or 1 2/5

OpenStudy (thecalchater):

and then 7/5+1/5= 8/5 or 1 3/5?

OpenStudy (freckles):

first sub-interval is [1,6/5] second sub-interval is [6/5,7/5] count to find the other 3 sub-intervals

OpenStudy (freckles):

yep the next right endpoint would be 8/5

OpenStudy (thecalchater):

[7/5.8/5] [8/5,9/5] and [9/5,10/5]

OpenStudy (freckles):

right and 10/5 is just 2 first sub-interval is [1,6/5] second sub-interval is [6/5,7/5] third sub-interval is [7/5,8/5] fourth sub-interval is [8/5,9/5] fifth sub-interval is [9/5,2]

OpenStudy (freckles):

so you take all of those left endpoints plug them into f(x)=1/x to find the height of each rectangle in the 2nd graph and you take all of those right endpoints plug them into f(x)=1/x to find the height of each rectangle in the 1st graph

OpenStudy (thecalchater):

ok do you want me to to that or do you want to?

OpenStudy (freckles):

you would probably be best to do it since this is your assignment :p

OpenStudy (thecalchater):

so then find the area of all of them and add them to get our answer?

OpenStudy (thecalchater):

when i plug it into my calc it says error...

OpenStudy (freckles):

the area of each rectangle is base*height where all the bases are equal so you can just add up all the heights and then multiply the sum by the base length of find the sum of all the areas

OpenStudy (thecalchater):

when i try to plug in for the heights it says error....

OpenStudy (thecalchater):

can you give me the decimals and ill multiply it by the base since my calc isnt doing it?

OpenStudy (freckles):

can't tell you why it says that 1/1 and 1/(6/5) and 1/(7/5) and 1/(8/5) and 1/(9/5) are all numbers

OpenStudy (thecalchater):

ohhhhh i was using 1/[7/5,7/6] lol give me a sec

OpenStudy (freckles):

that is the sub-intervals....

OpenStudy (freckles):

the list I just gave you was left endpoints of each of those sub-intervals we mentioned the right endpoints would exclude 1/1 and include 1/2 using that list I just mentioned

OpenStudy (thecalchater):

i got 1 5/7 5/8 5/9 so add them and multiply it by 2 to find the area?

OpenStudy (freckles):

what happened to 5/6 and also why are you multiplying by 2 2 is not the base length of each rectangle

OpenStudy (thecalchater):

its 1/5 for each

OpenStudy (freckles):

right

OpenStudy (thecalchater):

1879/504 for adding them together....... now multiply it by 1/5?

OpenStudy (thecalchater):

if so i got .74

OpenStudy (thecalchater):

0.7456349

OpenStudy (thecalchater):

that matches the upper sum of B. so it would be b?

OpenStudy (freckles):

\[\text{ area of rectangle }=\text{ base } \cdot \text{ height } \\ \\ \text{ the approximated area using } \\ \text{ \left endpoints } \\ \text{ recall the function is } f(x)=\frac{1}{x} \\ ... \\ \text{ so we have } \frac{1}{5} \cdot \frac{1}{1}+\frac{1}{5} \cdot \frac{1}{\frac{6}{5}}+\frac{1}{5} \cdot \frac{1}{\frac{7}{5}}+\frac{1}{5} \cdot \frac{1}{\frac{8}{5}}+\frac{1}{5} \cdot \frac{1}{\frac{9}{5}} \\ \text{ or } \frac{1}{5}(1+\frac{5}{6}+\frac{5}{7}+\frac{5}{8}+\frac{9}{5}) \approx .75\]

OpenStudy (freckles):

made a type-o 5/9 not 9/5

OpenStudy (freckles):

\[\text{ \right endpoint rule for this problem looks like } \\ \frac{1}{5}(\frac{5}{6}+\frac{5}{7}+\frac{5}{8}+\frac{5}{9}+\frac{1}{2})\]

OpenStudy (thecalchater):

so would it be b then?

OpenStudy (freckles):

if you want to verify it is b then evaluate the right endpoint rule thing I just wrote

OpenStudy (freckles):

but yep you should come out with that is b because the only choice that has the answer .75 is b

OpenStudy (thecalchater):

ok

OpenStudy (thecalchater):

can you check my other answers before we move to #3?

OpenStudy (thecalchater):

1D (we did) 2 b (we did) 3 4 5 B 6 A

OpenStudy (freckles):

5 and 6 appear to be right

OpenStudy (thecalchater):

ok now i just need help with 3 and 4 :)

OpenStudy (freckles):

this number 3 is similar to what we did just the graph isn't given this time

OpenStudy (freckles):

well and except it isn't finite amount of rectangles

OpenStudy (freckles):

and I suppose you could use a definite integral instead of using the whole summing thing...

OpenStudy (freckles):

the area between y=f(x) and y=0 on the interval from x=a to x=b can be found be evaluating \[\int\limits_a^b f(x) dx\]

OpenStudy (thecalchater):

here's the graph

OpenStudy (freckles):

dan already straight out gave the answer to number 5 earlier

OpenStudy (freckles):

oops number 4

OpenStudy (thecalchater):

oh i must not have seen it give me a sec

OpenStudy (freckles):

well we also said you find the height by using the...

OpenStudy (thecalchater):

left endpoints and sub integrals

OpenStudy (freckles):

not alone you have to plug them into the...

OpenStudy (thecalchater):

f(x) equation

OpenStudy (freckles):

yes c_i represents the sample point we are choosing in the sub-interval so f(c_i) represents the height we are using for each sub-interval

OpenStudy (thecalchater):

this is for #3 correct?

OpenStudy (freckles):

which ever one was about determining which one is the height

OpenStudy (thecalchater):

oh ok that's #4

OpenStudy (thecalchater):

oh i get it now.... ok last one number 3 finding the area of the bonded region

OpenStudy (freckles):

have you used what I said for number 3 yet? or did you want to find the answer using the proper way? the proper was being the way they want you to find it?

OpenStudy (thecalchater):

i posted the graph where do i go from there?

OpenStudy (thecalchater):

if you want post the steps out and ill follow along according to the other problem and find the final answer.....

OpenStudy (freckles):

ok we will go with the limits thing intead of the definite integral I suggested above \[\Delta x \cdot f(a+1 \Delta x)+\Delta x \cdot f(a+2 \Delta x)+\Delta x \cdot f(a+3 \Delta x)+ \cdots \Delta x \cdot f(a+n \Delta x) \\ =\Delta x \cdot [ f(a+1 \Delta x)+f(2+\Delta x)+ f(3+ \Delta x)+ \cdots +f(a+n \Delta x)] \\ =\Delta x \cdot \sum_{i=1}^n f(a+i \Delta x) \\ \\ \text{ so } \int\limits_a^b f(x) dx=\lim_{x \rightarrow \infty} \Delta x \cdot \sum_{i=1}^{n } f(x+i \Delta x)=Area\]

OpenStudy (freckles):

the limit way is the way they want you to go about it

OpenStudy (freckles):

the integral is another way you go about it if you had a choice

OpenStudy (thecalchater):

is this still for #4? I am talking about number 3. I understand #4

OpenStudy (freckles):

we are on number 3 number 4 was a short answer

OpenStudy (thecalchater):

ok want we me open a new question with that one and tag u?

OpenStudy (freckles):

well I kinda wrote everything you need for number 3 in this question ...

OpenStudy (thecalchater):

oh ok nvm then we can continue.